Please show mw how to solve for r in the given equation. A=P(1+r)(1+r) if A=2420 and P=2000
Answer by TimothyLamb(4379)

A = P(1+r)(1+r) if A=2420 and P=2000
---
A = P(1+r)(1+r)
A = P(rr + 2r + 1)
A = Prr + P2r + P
2420 = 2000rr + 4000r + 2000
2000rr + 4000r - 420 = 0
---
the above quadratic equation is in standard form, with a=2000, b=4000, and c=-420
---
to solve the quadratic equation, plug this:
2000 4000 -420
into this: https://sooeet.com/math/quadratic-equation-solver.php
---
the roots (solutions) of the quadratic are:
r = 0.1
r = -2.1
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php